### v1

@@ 2.8 Non-verbal Behaviour Generation
To enable situated interactions with multiple users at the same time, it is required that the non-verbal behaviour system components interface the social perception signals (presented in Sections 2.3, 2.4, 2.5 and 2.6) with the multi-party conversational manager (presented in Section 2.7). Among the set of social decisions that are required to be made by the non-verbal behaviour system we have: detect people’s arrival and departure, determines a person in the scene that wants or requires the robot’s attention, decide when to go, start an approach or guidance action, decide who to look at, switch the focus of attention during multi-party interactions, etc.
@@
The main functionality is to manage the currently available and running plans and to provide the functionality to send and receive information between different plans and the social state representations provided by the social scene understanding components. Through the interface with the social perception components (defined in the previous sections) the behaviour manager populates and maintains the planner’s knowledge base with information about the interaction, the social state, the people engaged in the interaction/conversation with the robot, etc. The behaviour manager module interfaces with the non-verbal behaviour generation through the ROS actions servers that control low-level action execution (gestures, navigation) for a pertinent social interaction.
@@
The behaviour generator provides mainly actions and behaviours for two aspects: First, it controls ARI’s arms, head, and eyes to generate gestures such as waving or pointing, and its gaze. The gesture and gaze controllers are hard-coded behaviours that can be called during dialogues.
Second, it provides a human-aware navigation controller to move among humans and join them for interactions. The controller faces two crucial challenges. Foremost, it needs to move safely and reliably due to its application around vulnerable persons. Secondly, it has to adhere to social norms regarding groups, for instance, respecting conversational groups and not moving through them, for example. In difference to existing methods [70, 71, 72], we opted to combine a Model Predictive Controller [73] with an explicit social space model of humans and groups [74], therefore called Social Navigation Controller (SNC) [75]. The SNC allows planning ARI’s trajectory precisely over a future time horizon while adhering to constraints. This provides a higher level of safety compared to other planning methods such as Dynamic Window Approaches [76] or learned end-to-end controllers such as by Reinforcement Learning [77]. Additionally, having an explicit social space model allows us to understand and tune it reliably compared to learned black-box models.
@@ 2.9 Experimenter Interface
During the experimentation, the robot will not be accompanied by computer science researchers or engineers, but by medical researchers and personnel. In order to conduct the experimental session, an appropriate experimenter-interface was developed. Through a conception-trial-update process, we converged on an interface design that is a trade-off between usability and controllability. The interface (Figure 6) is implemented for use on a dedicated tablet.
Figure 6: Snapshot of the Experimenter Interface.
@@ 3 Experiments
This section describes the experiments to validate the acceptability and usability of the introduced multi-modal conversational system on the ARI robot. The validation was conducted with real patients and companions at the Broca gerontology day-care hospital. The experimental protocol, measures, procedure, obtained results, as well as associated discussion and failure cases are described.
@@ 3.1 Experimental Protocol
This research was fully supported by the H2020 SPRING Project funded by the European Comission. As such, there are no conflicts of interest to be disclosed. As per the research involving humans and the informed consent, we provide details about the sought Ethical Committees and obtained Ethical Approvals in the following.
[…]
@@
Material
The experiments were conducted using the ARI robotic platform and the software architecture described in Section 2, as well as a dedicated external computing server. The external computing server has one NVIDIA RTX A6000 GPU with 48 GB of VRAM. Additionally, a dedicated secured 4G mobile connection is used for internet connection with a remote server running the LLM of the conversational system (Section 2.7). This remote server has 4x NVIDIA GeForce RTX 2080 Ti GPU. The collected experimental data is securely stored on a NAS server from Synology. An experimenter tablet was available to monitor the status of the robot, using the described experimental interface (Section 2.9), as well as for the experimenter to stop the robot and interaction if needed. An external camera was used to record the overall scene for posterior annotation and understanding.
@@ 3.2 Performance Measures
In this study, the user experience was assessed by two standardised scales. The first quantitative measure is the acceptability of the robot assessed with the Acceptability E-scale (AES) [81]. It is designed to measure the subjective acceptability of a system using 6 items, resulting in a global score that ranges from 6 to 30. For consumer ready-to-use products, the acceptability cut-off score is 25/30. The second quantitative measure is the usability of the robot assessed with the System Usability Scale (SUS) [82]. It is a 10-items scale designed to assess the overall user-friendliness of a system, and generating an overall score out of 100, where a higher score indicates better user-friendliness. For consumer ready-to-use products, the usability cut-off score is 72/100. Statistical differences were assessed using independent Student t-tests
@@ 3.3 Procedure
On the day of the test, the researcher checked that the tools (robot, tablet, camera) were working properly. Once the participants arrived, the researcher reminded them of the objectives of the study and the evaluation procedures. Participants could be escorted by one accompanying person. In case they confirmed their agreement to take part in the experiment, they were given a consent form to sign. The participant(s) were then invited to stand in front of the robot to begin the interaction with it. Two researchers led the activity, a facilitator, and a person who controlled the tablet and completed the observation grid. The facilitator’s role was to direct the test session and provide assistance or additional information to participants when needed.
@@ 3.4 Experimental Waves
Two separate waves of experimentation were conducted during this study, depending on the status of the architecture. The First wave (May 2023 – July 2023) used the initial version of the robot dialogue module, with an off-the-shelf automatic speech recognition model, and with almost no multi-party capabilities. In the Second wave (Sept. 2023 – Jan. 2024), some improvements to the system were made: The dialogue management included large-language models, the ASR was fine-tuned, the experimenter interface was updated, all-in-all to improve the user experience. The improvements were motivated by the participants’ feedback from the First wave.
@@ 3.5 Results
The recruitment efforts previous to the two waves lead to two sets of participants with slightly different age profiles for the patients, and wider age differences for the companions (see Table 1). The study counts a total of 63 participants with an overall average age of 74.5 (±12.31) years.
We also report the average and standard deviation of the AES and SUS scores (Table 2), split by wave and user group, as well as the overall score per wave.
3.6 Discussion
By looking at the main results, when we report the average AES and SUS results (Table 2), we observe a clear improvement between the First and Second wave, independently of the metric and for both the patient and companion group. A significant difference was found for the AES ($t(20)$$=$$-2.991$, $p$$<$.05.05.05.05) as the scores increased between the first ($M$$=$$15.5$, SD$=$$5.88$) and the second wave ($M$$=$$20.8$, SD$=$$5.20$). This significant difference was only found in patients ($t(18)$$=$$-2.906$, $p$$<$.05.05.05.05) and not in companions. However, no significant differences were observed for the SUS between the first and the second wave of experiments, both in patients and in companions. While there is an overall clear trend of improvement, the trend for the companion group is less pronounced than for the patient group.
Metric User Group First wave Second wave
AES Patients 14.7 (±05.73) 20.7 (±06.25)
Companions 18.0 (±04.64) 21.0 (±03.30)
Overall 15.5 (±05.88) 20.8 (±05.20)
SUS Patients 45.5 (±20.21) 56.8 (±12.63)
Companions 55.0 (±28.28) 57.5 (±10.40)
Overall 47.9 (±24.18) 57.0 (±22.88)
@@ 3.6 Discussion
We explain this improvement in acceptability and usability by the technical improvements we made to the software architecture (since the robot’s appearance did not change between waves): namely, the fine-tuning of the ASR module and the inclusion of LLMs in the conversation manager. These two major modifications, together with bug fixes, allowed a more natural interaction with the users since (i) ARI understood better what the participants were saying and (ii) ARI was able to answer to a wider range of questions with reasonable (although now always exact) answers. In particular, the participant’s feedback reported: enjoyment in using the robot, usefulness, acceptability of the robot’s reaction time (talking) and overall satisfaction with the robot.

### v2

@@ 2.8 Non-verbal Behaviour Generation
To enable situated interactions, meaning interactions anchored in the physical and social environments of the robot, with multiple users at the same time, it is required that the non-verbal behaviour system components interface the social perception signals (presented in Sections 2.3, 2.4, 2.5 and 2.6) with the multi-party conversational manager (presented in Section 2.7). Among the set of social decisions that are required to be made by the non-verbal behaviour system, we have: detect people’s arrival and departure, determines a person in the scene that wants or requires the robot’s attention, decide when to go, start an approach or guidance action, decide who to look at, switch the focus of attention during multi-party interactions, etc.
@@
The main functionality is to manage the currently available and running plans and to provide the functionality to send and receive information between different plans and the social state representations provided by the social scene understanding components. Through the interface with the social perception components (defined in the previous sections), the behaviour manager populates and maintains the planner’s knowledge base with information about the interaction, the social state, the people engaged in the interaction/conversation with the robot, etc. The behaviour manager module interfaces with the non-verbal behaviour generation through the ROS actions servers that control low-level action execution (gestures, navigation) for a pertinent social interaction.
@@
The behaviour generator provides mainly actions and behaviours for two aspects: First, it controls ARI’s arms, head, and eyes to generate gestures such as waving or pointing and its gaze. The gesture and gaze controllers are hard-coded behaviours that can be called during dialogues.
Figure 6: Snapshot of the Experimenter Interface.
Second, it provides a human-aware navigation controller to move among humans and join them for interactions. The controller faces two crucial challenges. Foremost, it needs to move safely and reliably due to its application around vulnerable persons. Secondly, it has to adhere to social norms regarding groups, for instance, respecting conversational groups and not moving through them, for example. In difference to existing methods [78, 79, 80], we opted to combine a Model Predictive Controller [81] with an explicit social space model of humans and groups [82], therefore called Social Navigation Controller (SNC) [83]. The SNC allows planning ARI’s trajectory precisely over a future time horizon while adhering to constraints. This provides a higher level of safety compared to other planning methods such as Dynamic-Window Approaches [84] or learned end-to-end controllers such as by Reinforcement Learning [85], see [25] for a broader perspective. Additionally, having an explicit social space model allows us to understand and tune it reliably compared to learned black-box models.
@@ 2.9 Experimenter Interface
During the experimentation, the robot will not be accompanied by computer science researchers or engineers, but by medical researchers and personnel. In order to conduct the experimental session, an appropriate experimenter interface was developed. Through a conception-trial-update process, we converged on an interface design that is a trade-off between usability and controllability. The interface (Figure 6) is implemented for use on a dedicated tablet.
@@ 3 Experiments
This section presents the experiments to validate the acceptability and usability of the introduced multi-modal conversational system on the ARI robot. The validation was conducted with real patients and companions at the Broca gerontology day-care hospital. The experimental protocol, measures, procedure, obtained results, as well as associated discussion and failure cases, are described in the following.
@@ 3.1 Experimental Protocol
This research was fully supported by the H2020 SPRING Project funded by the European Commission. As such, there are no conflicts of interest to be disclosed. As per the research involving humans and the informed consent form, we provide details about the sought Ethical Committees and obtained Ethical Approvals in the following.
The study was approved by the French National Ethical Review Board Comité National de Protection des Personnes, CPP Ouest II, Maison de la Recherche Clinique-CHU Angers (approval number: 2021/20), the local ethics committee of the University of Paris (Comité d’Ethique de la Recherche CER-N IRB: 00012020-108), and was compliant with the General Data Protection Regulation (DPO: 20210114153645 register AP-HP). The study is also registered as a clinical trial. [^8] https://clinicaltrials.gov/study/NCT05089799 Additional information can be found in [88]. In order to guarantee universal healthcare access, the Ethical Approval restricted the experiments in the hospital to an auxiliary room, instead of the main waiting room, in a way that patients unwilling to interact with the robot would not feel aside or unwelcome. This implies that only the participants who signed the informed consent were seen and heard by the robot, and data from no other person is included in the recordings or experiments. The auxiliary room is small ($<20$ m2), making the robot self-localisation and navigation tasks trivial. Hence, both skills have been validated in our respective laboratories and when possible in the Broca hospital without patients.
@@
Summarising, for this study involving an experimental product, we took care of respecting the rights and dignity of users through two main measures:
  • Informed consent: Before taking part in the experiments, all users were asked to sign consent forms and image release forms.
  • Anonymity and data protection: The experiments took place in a closed room, which ensured that no participant who had not signed the consent form was filmed. All the videos recorded by the robot and any non-anonymous data were carefully processed and deleted, thus protecting users’ privacy. This is a manual process done by the experimenter, previous to sharing the data with any of the project’s partners.
Computational and Storage Resources
The experiments were conducted using the ARI robotic platform and the software architecture described in Section 2, as well as a dedicated external computing server. The external computing server has one NVIDIA RTX A6000 GPU with 48 GB of VRAM. Additionally, a dedicated secured 4G mobile connection is used for an internet connection with a remote server running the LLM of the conversational system (Section 2.7). This remote server has four NVIDIA GeForce RTX 2080 Ti GPU. Summarising, the computations involving heavy data flows (images, audio) are run on an external computer connected to ARI via Wi-Fi, while the dialogue, requiring only textual information, is run on the server of one of the project’s partners. The overall computing pipeline is privacy-preserving while ensuring enough reactivity on ARI’s side. The collected experimental data is securely stored on a NAS server from Synology, after pseudonymisation. The server was set up following the instructions of Inria’s Chief Security Officer (in a locked closet, encrypted, and password protected). An experimenter tablet was available to monitor the status of the robot, using the described experimental interface (Section 2.9), as well as for the experimenter to stop the robot and interaction if needed. An external camera was used to record the overall scene for posterior annotation and understanding.
@@ 3.2 Performance Measures
In this study, the user experience was assessed by two standardised scales. The first quantitative measure is the acceptability of the robot assessed with the Acceptability E-scale (AES) [89]. It is designed to measure the subjective acceptability of a system using 6 items, resulting in a global score that ranges from 6 to 30. For consumer ready-to-use products, the acceptability cut-off score is 25/30. The second quantitative measure is the usability of the robot assessed with the System Usability Scale (SUS) [90]. It is a 10-items scale designed to assess the overall user-friendliness of a system, and generate an overall score out of 100, where a higher score indicates better user-friendliness. For consumer ready-to-use products, the usability cut-off score is 72/100. Statistical differences were assessed using independent Student t-tests. The details of the AES and SUS scales can be found in Appendix A. In addition to these two quantitative measures, we have also conducted a semi-structured interview on ethical issues, containing six open questions, ranging from the robot design to the potential emotional attachment. Other strategies could have consisted in using the story dialogue method [91]. The questions and sample answers will be presented in the results section.
@@ 3.3 Procedure
On the day of the test, the researcher checked that the tools (robot, tablet, camera) were working properly. Once the participants arrived, the researcher reminded them of the objectives of the study and the evaluation procedures. Participants could be escorted by one accompanying person that could participate in the experiment together with the patient. In some cases, accompanying adults were also invited to take part in the experiment on their own during their companions’ appointments. In case they confirmed their agreement to take part in the experiment, every participant was given a consent form to sign. The participant(s) were then invited to stand in front of the robot to begin the interaction with it in French. Two researchers led the activity, a facilitator, and a person who controlled the tablet and completed the observation grid. The facilitator’s role was to direct the test session and provide assistance or additional information to participants when needed. Overall, the entire interaction with one participant lasted between five and ten minutes, depending on the willingness of the participant to pursue discussing with ARI.
Figure 7: Experimental set up at the Broca day-care gerontology hospital. Two experimenters monitor the interaction of a patient with the ARI robot. Everything is recorded both by ARI’s sensors and by the external camera.
The experiments were documented using several devices. The robot’s onboard cameras and microphones captured the interactions from an ego-centric point of view, while an external camera (1 meter away from the robot) filmed the entire scene from an external perspective for further analysis. During the interaction with the robot, the experimenter took photos of the exchange via a laboratory camera. The interviews were recorded with a dictaphone by the experimenter, who then manually transcribed the conversations, ensuring a faithful and accurate transcription for qualitative data analysis. For all the experiments, we obtained $1,142$ minutes of interviews excluding interaction with the robot, representing $418$ pages of transcriptions. The data was then analyzed by three laboratory researchers, who dematerialized the interviews using the Dedoose [^11] https://www.dedoose.com/ software. Figure 7 shows the experimental setup in the day-care hospital.
@@ 3.4 Experimental Waves
Two separate waves of experimentation were conducted during this study, depending on the status of the architecture. The experiments within the same wave used the exact same software architecture, while experiments in different waves correspond to changes in some of the modules. The architecture hence evolved following the feedback the users provided, which was exploited to improve certain software modules as detailed below. The first wave ($1^{\text{st}}$: May 2023 – July 2023) used the initial version of the robot dialogue module, with an off-the-shelf automatic speech recognition model, and with almost no multi-party capabilities. In the second wave ($2^{\text{nd}}$: Sept. 2023 – Jan. 2024), some improvements to the system were made: The dialogue management included large-language models, the ASR was fine-tuned, the experimenter interface was updated, all-in-all to improve the user experience.
@@ 3.5 Results
The recruitment efforts previous to each experimental wave lead to three sets of participants with slightly different age profiles for the patients, and wider age differences for the companions (see Table 1). The study counts a total of 63 participants with an overall average age of 74.5 (±12.31) years. We also report the average and standard deviation of the AES and SUS scores (Table 2), split by wave and user group, as well as the overall score per wave. In addition, we have computed the Cronbach’s alpha coefficient for the two scales, resulting in $\alpha=0.87$ and $\alpha=0.86$ for the AES and SUS scales respectively, thus indicating internal consistency of the two measurement scales.
Wave User Group AES SUS
$1^{\text{st}}$ Patients 14.7 (±05.73) 45.5 (±20.21)
Companions 18.0 (±04.64) 55.0 (±28.28)
Overall 15.5 (±05.88) 47.9 (±24.18)
$2^{\text{nd}}$ Patients 20.7 (±06.25) 56.8 (±12.63)
Companions 21.0 (±03.30) 57.5 (±10.40)
Overall 20.8 (±05.20) 57.0 (±22.88)
@@
Table 3: Verbatim answers to the questions of the semi-structured interview on ethical concerns, structured by the experimental waves.
Q1. Do you approve of robots that look like humans?
“The design is a bit too mechanical, it lacks human warmth.”
“I like the futuristic style, it fits with my idea of robots.”
“The design is clean, but it could be more user-friendly.”
“It’s more fun and more attractive that the robot has a humanoid appearance.”
“The robot shouldn’t be too realistic. It should not be an identical replica of the human, especially if it is to be used in public places.”
Q2. What do you think of the use of social robots in hospitals?
“I don’t think it’s really suitable for hospitals.”
“It could be useful, but it needs to be better integrated into workflows.”
“I think this robot could really lighten some tasks.”
“It must not be to replace the human being. A robot that comes to help a caregiver lift a person who has fallen, I agree with that, but a robot that comes to do the consultation… no way.”
“The robot can be a precious aid, of help and assistance, but it should be used in moderation.”
Q3. Do you think it’s right/just for robots to record and store information?
“I’m not comfortable with the idea of my data being stored.”
“I’m not sure what the robot is doing with my information.”
“I’d like to know exactly who has access to my data.”
“If it’s collected online, no. If it’s protected, like with computers, it’s fine.”
“Data is anyway collected whether it is recorded by a robot or by a computer in the hospital.”
Q4. Are you concerned about the mistakes social robots could make?
“If the robot makes a mistake, it can have serious consequences, especially in a hospital environment.”
“It’s a bit worrying because a simple mistake could disorient someone.”
“The robot made a mistake, but it made up for it quite quickly. That’s reassuring.”
“It can happen indeed [that the robot makes mistakes], but like any software, or any computer system, we must not worry about everything, we must progress.”
“If the robot doesn’t provide the right direction, it causes delay and irritation. But it’s not like a medical intervention where it would be really dramatic.”
Q5. Do you think it’s possible to have an emotional bond with a robot?
“I think this robot could create a kind of emotional dependency.”
“I wonder if it’s healthy to get attached to a robot.”
“We could easily feel too close to the robot, that’s a worry.”
“The emotional bond can only develop from one living person to another living person. For there to be affection, there has to be a relationship, you have to be able to touch, you have to be able to feel.”
“I don’t think so, because for me it [the robot] is a machine.”
Q6. Are you concerned about social robots replacing human beings?
“It could replace humans in certain tasks.”
“It’s scary to think that this could replace humans.”
“I think some jobs will be threatened by this kind of technology.”
“From a social point of view, yes, the risk of replacement is a bit worrying. Robots are useful in industry, to accomplish repetitive tasks and why not in a hospital. Then the staff at the reception desk and other staff will be unemployed if they are replaced by robots.”
“Robots cannot replace people, they can only help. It’s not the same thing. A machine will never replace a human, it’s impossible.”
Finally, Table 3 reports que questions of the semi-structured ethical interview, and two verbatim answers for each question, at each experimental wave. These answers were provided by patients after having completed the interaction experiment with ARI. The verbatim answers were analyzed following an inductive thematic analysis [93]. Through careful reading, it consists of identifying, analysing, and reporting themes within the data, with the aim of capturing useful insights about the data in relation to our research objectives. We chose to use this method because it offers several advantages such as being very flexible (not bound to a specific theoretical framework compared to the deductive approach for instance) and being more data-oriented as the themes emerge directly from the data. Thus, we can avoid any possible bias such as the researcher’s possible preconceptions. Furthermore, as our experiments took place in real-life settings with a maturing system (a social robot within a hospital), we chose an open-minded approach in order not to miss anything important.
3.6 Discussion
[…]
We explain this improvement in acceptability and usability by the technical improvements we made to the software architecture (since the robot’s appearance did not change between waves): namely, the fine-tuning of the ASR module and the inclusion of LLMs in the conversation manager. These two major modifications, together with bug fixes, allowed a more natural interaction with the users since (i) ARI understood better what the participants were saying and (ii) ARI was able to answer to a wider range of questions with reasonable (although now always exact) answers. In particular, the participant’s feedback reported: enjoyment in using the robot, usefulness, acceptability of the robot’s reaction time (talking), and overall satisfaction with the robot.
@@ 3.6 Discussion
Regarding the semi-structured interview dealing with ethical issues (see verbatim excerpts in Table 3), we extract the following main messages summarising our experience. Generally speaking, participants found that ARI’s design was acceptable but could be improved, in the sense that it looked a bit too mechanical and would benefit from a warmer look. The robot’s usefulness was globally acknowledged, but it was also pointed out that there was room for improvement in terms of integration in the workflow (we did not target such integration in our experiments). Data use and privacy are major concerns, and participants expressed their need for transparency in terms of data use, sharing, storage, etc. We believe this awareness demonstrates that older adults are well informed about the dangers of misuse of personal data, and we interpret this request for transparency as a very good societal sign. We also reiterate that for our experiments the data was not shared beyond the SPRING consortium, in agreement with the Ethical Approvals and the protocol validated by Inria’s Chief Security Officer. The intertwining between the first and third questions (appearance and privacy) appears to be a complex matter [94], since users link certain appearance features to their privacy being or not preserved by the robot. Regarding the potential mistakes made by the robot, the participants were moderately concerned but felt reassured by having human supervision. Participants were also aware of the risk of attachment to the robot, and raised the concern of preventing this from happening, especially with vulnerable persons. Finally, the participants were significantly concerned about the potential replacement of human jobs by robots, pointing out that they would be useful for specific small tasks. Overall, these verbatim answers provided an interesting view of how older adults perceive the use of socially assistive robots in healthcare facilities, and they are very complementary to the quantitative acceptability and usability scales results discussed above. In addition, some of the ethical concerns expressed by the participants (loss of privacy, decreased human contact/human replacement) are aligned with existing ethical discussions in the literature [95], while others (e.g. developing an emotional atttachment) are complementary.
All these experiments were possible thanks to the software architecture described in Section 2 (as well as to the designed experimental protocol). To our knowledge, this is the first robotic software architecture enabling audio-visual interaction of a robotic platform with multiple persons. It is also one of the first architectures exploiting the ROS4HRI standard for human-robot interaction. Very importantly, this is the first time such an advanced interaction architecture has been validated with more than 60 experiments with real patients and companions in a day-care gerontological hospital.